# LoSiA: Efficient High-Rank Fine-Tuning via Subnet Localization and Optimization

### v3

@@ 5 Conclusion
We present LoSiA, a novel PEFT framework that dynamically identifies and optimizes core sub-networks. Through sensitivity-based localization, asynchronous re-selection, and efficient high-rank adaptation, LoSiA achieves high throughput and low activation overhead. Extensive experiments show that LoSiA outperforms baselines on domain-specific and common-sense reasoning tasks while reducing forgetting. We hope that our work will inspire future research to further explore the intrinsic substructures in supervised fine-tuning.
@@ A.1.1 Proof for Formula 1
On a batch $\mathcal{B}$ composed of $M$ samples, the MSE Loss between full fine-tuning (produces model $f$) and training with parameter set $P$ (produces model $f^{\prime}$) is given by:
@@
In SGD optimizer, suppose learning rate is $\eta$, the difference in fine-tuned parameter is:
$W-W^{\prime}=-\eta\mathbb{1}_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})$ (12)
@@
$\mathcal{L}_{MSE}$ $\leq\eta^{2}\frac{\|\mathbb{1}_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})\|^{2}_{F}\|x\|^{2}_{F}}{M}$ (13)
The result suggests that maximizing the sum of $\nabla_{W_{0}}\mathcal{L}(\mathcal{B})_{ij}$ where $(i,j)\in P$ ideally tightens the approximate error of training on subset.
@@
Similarly, since $W-W^{\prime}=-\eta\mathbb{1}_{(i,j)\not\in P}\cdot\tilde{G_{t}}$, we analysis relationship between $\tilde{G_{t}}$ and $G_{t}$:
@@
Suppose $M_{t}>0$, when $G_{t}<\frac{(1-\beta_{1})V_{t}}{(1-\beta_{2})M_{t}}$, $\frac{\partial(\tilde{G_{t}})^{2}}{\partial G_{t}}>0$. In practice, the typical settings are $\beta_{1}=0.9,\beta_{2}=0.999$. Therefore, when $G_{t}<10^{2}\frac{V_{t}}{M_{t}}$, $\tilde{G}_{t}$ increases with $G_{t}$, effectively covering a broad range of non-stationary optimization scenarios.
@@ A.1.2 Proof for Formula 3
The foundational work was established by LeCun et al. (1989) and Kirkpatrick et al. (2016). However, for real-time importance calculation during training, approximations is necessary and is derived bellow. Element-wise importance score $I(\cdot)$ is formulated as:
$\begin{split}I(W_{k})=&|\Delta\mathcal{L}(\mathcal{D})|=|\mathcal{L}(\mathcal{D})-\mathcal{L}_{W_{k}=0}(\mathcal{D})|\\ =&|\frac{\partial\mathcal{L}^{T}(\mathcal{D})}{\partial{W_{k}}}W_{k}-\frac{1}{2}W_{k}^{T}H_{kk}W_{k}\\ &+o(W_{k}^{2})|\end{split}$ (19)
where $H$ denotes the Hessian matrix. It is computational intensive for Hessian calculation, we therefore use fisher information matrix $F$ for diagonal elements instead:
@@
Approximating the expectation by dataset $\mathcal{D}$ based on Monte Carlo Method, it derives:
@@
During training, the dataset $\mathcal{D}$ is processed in batches $\mathcal{B}_{i}$, and the batch gradient is calculated as $\nabla_{W}\mathcal{L}(\mathcal{B}_{i})=\frac{1}{M}\sum_{j=1}^{M}\nabla_{W}\mathcal{L}(\mathcal{B}_{ij})$. To avoid calculating the gradients separately for each sample in the batch, we approximate $\sum_{(x,y)\in\mathcal{B}_{i}}\frac{1}{M}(\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}})^{2}$ to the term $(\frac{\sum_{(x,y)\in\mathcal{B}_{i}}\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}}{M})^{2}$. To analyze errors, assume $g=\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}\sim G$, we have:
@@ A.1.3 Maximizing Formula 7 is NP-Hard
Given an arbitrary matrix $A^{n\times m}$, select $\tilde{n}\ (\tilde{n}\leq n)$ rows $X_{S}$ and $\tilde{m}\ (\tilde{m}\leq m)$ columns $Y_{S}$, to maximize the sum $\sum_{i\in X_{S}}\sum_{j\in Y_{S}}A_{ij}$.
@@
Construct a special form of $A^{n\times n}$ as the adjacent matrix of graph $G$ with larger values on diagonal maximum, that is:
@@
Then, MCP can be reduced to the task in polynomial time following the algorithm:
@@
Therefore, a NP-Complete problem can be reduced to the task in polynomial time, which yields the conclusion that the task is NP-Hard.
@@ A.2.1 Gradient Magnitude Distribution
To investigate the universality of the sparse sub-network structure for large gradients, we analyze gradient magnitude distributions across different layers, as shown in Figure 9. Both Gradients of the Self-Attention and MLP modules exhibits the consistent structure of core subnets.
@@ A.3.2 Common-Sense Reasoning Tasks
The datasets of common-sense reasoning tasks are presented in Table 8, while corresponding hyperparameters detailed in Table 9. The GPU memory usage remain aligned. For each PEFT baselines, searches in learning rate are performed.
@@ A.3.3 Rank Factor Robustness
To evaluate the impact of the rank factor $p$, which determines the scale of the core subnets, we conduct an ablation study on MetaMathQA. The results demonstrate LoSiA’s robustness across various subnet scales. Note that $p=1/16$ may be relatively small for effective subnet fine-tuning, while increasing computational budget boosts the performance.
Table 10: Rank Factor Robustness on GSM8K
@@ A.3.4 Continue Learning
To examine whether reduction of intruder dimensions in LoSiA mitigates forgetting in continue learning, we sequentially adapt LLaMA-2 7B through five common-sense reasoning tasks by the order HellaSwag, PIQA, BoolQ, SIQA and Winogrande. Learning rate for LoRA is 1e41e-4 and for LoSiA is 5e55e-5. The remaining hyperparameters are consistent with Table 9. LoRA modules are merged into the backbone before subsequent task adaptation.
Suppose the model learn sequentially on $N$ tasks. Let $P_{i,j}$ denote the ACC on task $j$ after training on task $i$. Following Zhao et al. (2024b), we formulate the metrics (AP, FWT and BWT) as bellow:
Average Performance: $AP=\frac{1}{N}\sum_{i=1}^{N}P_{N,i}$
Backward Transfer: The metric evaluates the impact of learning later tasks on the model’s performance on an earlier task, that is $\text{BWT}=\frac{1}{N-1}\sum_{i=1}^{N-1}(P_{N,i}-P_{i,i})$.
@@
[…]
@@ A.4.1 Memory Estimate
Consider a model with $L$ decoder layers, each containing $K$ tunable matrices. The model use $b$-bit precision storage, with hidden dimension $d$ and vocabulary size $V$. Table 13 shows GPU memory consumption details of LoRA, GaLore and LoSiA.
Table 13: Comparison of Memory Consumptions. Cells in green highlight the components that may notably lower than other methods, while in red highlight the components that may cause relatively large memory consumption.
@@
For optimizers like AdamW, LoSiA reduces the gradient dimension of the output layer to a fraction $p_{o}$, while GaLore performs full fine-tuning on the output layer of shape $d\times V$. Both GaLore and LoSiA utilize per-layer weight update techniques for gradient computation. In contrast, basic LoRA implementations require collecting all gradients after backward propagation. The reduced memory overhead of gradient storage allows GaLore and LoSiA to fully train the $\mathtt{lm\_head}$ in LLaMA-2 models.
In terms of auxiliary parameters, GaLore requires storing down- and up-projection matrices. Since $R$ is typically high-rank, GaLore’s auxiliary parameters can be significantly larger than those of other methods.
Regarding total memory consumption, increasing the rank in LoRA and GaLore incurs substantial overhead. However, in LoSiA, only the term $2(LKd^{2}p^{2}b+Vdp_{o}b)$ scales with rank factor $p$, resulting in a more efficient memory footprint.
Table 14: Details of Trainable Parameters for LoRA and LoSiA under Different Hyperparameter Configurations on LLaMA-2 7B.
LoRA
$\mathbf{Rank\ r}$ $16$ $64$ $256$ $1024$
$\mathbf{\#Trainable}$ 40.040.0M 160.0160.0M 640.0640.0M 2560.02560.0M
$\mathbf{Mem(GB)}$ $22.33$ $23.72$ $28.42$ $63.90$
@@ A.4.2 Latency Measurement
While demonstrating superior performance among existing baselines, LoSiA reduces training latency by 14.1% compared to LoRA, 55.8% compared to DoRA. The acceleration is mainly due to the elimination of low-rank matrix multiplication. Specifically, during backward propagation with Gradient Checkpointing, the production of low-rank matrices introduces significant overhead for activation recomputation and gradient calculation. Note that LoRA can avoid gradient calculations on backbone weights, but this requires specialized implementations and introduces a large coefficient of computational complexity.

### v4

@@ 5 Conclusion
We present LoSiA, a novel PEFT framework that dynamically identifies and optimizes core sub-networks. Through sensitivity-based localization, asynchronous re-selection, and efficient high-rank adaptation, LoSiA achieves high throughput and low activation overhead. Extensive experiments show that LoSiA outperforms baselines on domain-specific and common-sense reasoning tasks while reducing forgetting. We hope that our work will inspire future research to further explore intrinsic substructures in supervised fine-tuning.
@@ 6 Limitation
Acknowledgements
This work is supported by the National Natural Science Foundation of China (No. 52539001). It’s partially supported by CHN Energy Dadu River Big Data Services Co., Ltd. It’s also supported by Student Research Training (SRT) project of Tsinghua University. We also thank anonymous reviewers for their valuable feedback.
@@ A.1.1 Proof for Formula 1
On a batch $\mathcal{B}$ composed of $M$ samples, the MSE loss between full fine-tuning (which produces model $f$) and training on parameter set $P$ (which produces model $f^{\prime}$) is given by:
@@
In SGD optimizer, supposing the learning rate is $\eta$, the difference in fine-tuned parameter is:
$W-W^{\prime}=-\eta 1_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})$ (12)
@@
$\mathcal{L}_{MSE}$ $\leq\eta^{2}\frac{\|1_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})\|^{2}_{F}\|x\|^{2}_{F}}{M}$ (13)
The result suggests that maximizing the sum of $\nabla_{W_{0}}\mathcal{L}(\mathcal{B})_{ij}$ where $(i,j)\in P$ ideally tightens the approximate error of training on parameter subset.
@@
Similarly, since $W-W^{\prime}=-\eta 1_{(i,j)\not\in P}\cdot\tilde{G_{t}}$, we analyze the relationship between $\tilde{G_{t}}$ and the original gradient $G_{t}$ by element:
@@
Suppose $M_{t}>0$, when $G_{t}<\frac{(1-\beta_{1})V_{t}}{(1-\beta_{2})M_{t}}$, $\frac{\partial(\tilde{G_{t}})^{2}}{\partial G_{t}}>0$. In practice, typical settings are $\beta_{1}=0.9,\beta_{2}=0.999$. Therefore, when $G_{t}<10^{2}\frac{V_{t}}{M_{t}}$, $\tilde{G}_{t}$ increases with $G_{t}$, effectively covering a broad range of non-stationary optimization scenarios.
@@ A.1.2 Proof for Formula 3
The foundational work was established by LeCun et al. (1989) and Kirkpatrick et al. (2016). However, to establish real-time importance calculation during training, approximations are necessary and are derived below. Element-wise importance score $I(\cdot)$ is formulated as:
$\begin{split}I(W_{k})=&|\Delta\mathcal{L}(\mathcal{D})|=|\mathcal{L}(\mathcal{D})-\mathcal{L}_{W_{k}=0}(\mathcal{D})|\\ =&|\frac{\partial\mathcal{L}^{T}(\mathcal{D})}{\partial{W_{k}}}W_{k}-\frac{1}{2}W_{k}H_{kk}W_{k}\\ &+o(W_{k}^{2})|\end{split}$ (19)
where $H$ denotes the Hessian matrix, which is computationally intensive. Therefore, the fisher information matrix $F$ is used instead to obtain diagonal elements of the Hessian matrix:
@@
Approximating mathematical expectations on dataset $\mathcal{D}$ using the Monte Carlo method derives:
@@
During training, the dataset $\mathcal{D}$ is processed in batches $\mathcal{B}_{i}$, and the batch gradient is calculated as $\nabla_{W}\mathcal{L}(\mathcal{B}_{i})=\frac{1}{M}\sum_{j=1}^{M}\nabla_{W}\mathcal{L}(\mathcal{B}_{ij})$. To avoid calculating the gradients separately for each sample in the batch, we approximate $\sum_{(x,y)\in\mathcal{B}_{i}}\frac{1}{M}(\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}})^{2}$ to the term of $(\frac{\sum_{(x,y)\in\mathcal{B}_{i}}\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}}{M})^{2}$. To analyze errors, assume $g=\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}\sim G$, we have:
@@ A.1.3 Maximizing Formula 7 is NP-Hard
Given an arbitrary non negative matrix $A^{n\times m}$ and cardinal budget requirements $\tilde{n},\tilde{m}$. Select $\tilde{n}$ rows $X_{S}$ and $\tilde{m}$ columns $Y_{S}$ to maximize the sum $\sum_{i\in X_{S}}\sum_{j\in Y_{S}}A_{ij}$.
@@
Construct a special form of $A^{n\times n}$ as the adjacent matrix of graph $G$ with larger values on the diagonal maximum, that is:
@@
Then, the MCP problem can be reduced to the task in polynomial time following the algorithm below:
@@
Therefore, an NP-Complete problem can be reduced to the task in polynomial time, which yields the conclusion that the task is NP-Hard.
@@ A.2.1 Gradient Magnitude Distribution
To investigate the universality of the sparse sub-network structure for large gradients, we analyze gradient magnitude distributions across different layers, as shown in Figure 9. Both Gradients of the Self-Attention and MLP modules exhibit the consistent structure of core subnets. We also quantify this observation on different layers and each submodule in Table 6. Core-subnet localization markedly outperforms random selection and closely approaches the ideal Top-K baseline, corroborating the validity of sparse-gradient patterns. Note, however, that the ideal Top-K set is irregular and therefore entails nontrivial runtime overhead.
Table 6: Sum of Absolute Gradient Values ($\times 10^{3}$) for Selection Patterns. "Subnet" stands for the core subnet localization algorithm, while Top-K is an ideal selection.
Layer Submodule Total Random Subnet Top-K (Ideal)
5 q_proj $16.38$ $1.02$ $3.82$ $6.72$
k_proj $14.02$ $0.88$ $3.18$ $5.70$
v_proj $83.97$ $5.25$ $18.69$ $30.59$
o_proj $93.18$ $5.82$ $19.20$ $29.70$
up_proj $133.12$ $8.32$ $20.48$ $33.02$
down_proj $144.38$ $9.02$ $21.89$ $37.12$
gate_proj $101.89$ $6.37$ $16.51$ $27.26$
15 q_proj $15.49$ $0.97$ $4.83$ $6.62$
k_proj $12.29$ $0.77$ $3.22$ $5.25$
v_proj $48.64$ $3.04$ $12.93$ $16.90$
o_proj $48.38$ $3.02$ $12.29$ $14.98$
up_proj $65.02$ $4.06$ $15.30$ $20.86$
down_proj $70.14$ $4.38$ $15.42$ $21.63$
gate_proj $54.53$ $3.41$ $13.63$ $19.07$
25 q_proj $6.94$ $0.43$ $2.82$ $4.13$
k_proj $6.02$ $0.38$ $1.90$ $3.54$
v_proj $12.93$ $0.81$ $4.10$ $5.98$
o_proj $13.38$ $0.84$ $4.61$ $5.86$
up_proj $32.51$ $2.03$ $8.13$ $10.82$
down_proj $36.10$ $2.26$ $8.64$ $11.20$
gate_proj $25.09$ $1.57$ $6.50$ $8.70$
@@ A.3.2 Common-Sense Reasoning Tasks
The datasets of common-sense reasoning tasks are presented in Table 9, while corresponding hyperparameters detailed in Table 10. The GPU memory usage remains aligned. For each PEFT baselines, searches in learning rate are performed.
@@ A.3.3 Rank Factor Robustness
To evaluate the impact of the rank factor $p$, which determines the scale of selected core subnets, we conduct an ablation study on MetaMathQA as Table 11 shows. Performance grows steadily with the number of training parameters on both Gemma 2B and LLaMA-2 7B. The results demonstrate LoSiA’s robustness across various subnet scales. Note that $p=1/16$ may be relatively small for effective subnet fine-tuning, while increasing the computational budget boosts the performance.
Table 11: Rank Factor Robustness of LoSiA on GSM8K
@@ A.3.4 Continue Learning
To examine whether reduction of intruder dimensions in LoSiA mitigates forgetting in continue learning, we sequentially adapt LLaMA-2 7B through five common-sense reasoning tasks: HellaSwag, PIQA, BoolQ, SIQA and Winogrande. Learning rate for LoRA is 1e41e-4 and for LoSiA is 5e55e-5. The remaining hyperparameters are consistent with Table 10. LoRA modules are merged into the backbone before subsequent task adaptation.
@@
Method Task (#1) HellaS (#2) PIQA (#3) BoolQ (#4) SIQA (#5) WinoG ST
Seq-LoRA HellaSwag $59.86$ $55.64$ $59.10$ $57.86$ $54.36$ $59.86$
PIQA $76.01$ $80.52$ $77.86$ $78.73$ $77.64$ $79.33$
BoolQ $77.80$ $73.27$ $86.30$ $80.12$ $75.93$ $88.07$
SIQA $45.80$ $47.80$ $45.85$ $59.52$ $46.11$ $56.86$
Winogrande $64.25$ $68.35$ $68.82$ $69.93$ $79.08$ $73.88$
Seq-LoSiA HellaSwag $63.72$ $61.89$ $61.11$ $60.37$ $\mathbf{56.43}$ $63.72$
PIQA $78.29$ $79.49$ $79.82$ $79.38$ $\mathbf{77.75}$ $81.50$
BoolQ $77.52$ $70.76$ $83.24$ $82.54$ $\mathbf{81.99}$ $84.13$
SIQA $47.80$ $48.26$ $48.26$ $59.93$ $\mathbf{56.04}$ $61.05$
Winogrande $68.51$ $67.88$ $68.51$ $71.82$ $\mathbf{80.19}$ $77.19$
@@ A.4.1 Memory Estimate
Consider a model with $L$ decoder layers, each containing $K$ tunable matrices. The model use $b$-bit precision storage, with hidden dimension $d$ and vocabulary size $V$. Table 14 shows GPU memory consumption details of LoRA, GaLore and LoSiA. For optimizers like AdamW, LoSiA reduces the gradient dimension of the output layer to a fraction $p_{o}$, while GaLore performs full fine-tuning on the output layer of shape $d\times V$. Both GaLore and LoSiA utilize per-layer weight update techniques for gradient computation. The update is therefore computed upon gradient acquisition and then promptly discarded.
In terms of auxiliary parameters, GaLore requires storing down- and up-projection matrices. Since $R$ is typically high-rank, GaLore’s auxiliary parameters can be significantly larger than those of other methods, which may induce additional GPU memory consumptions.
Regarding total memory consumption, increasing the rank in LoRA and GaLore incurs substantial overhead. However, in LoSiA, only the term $2(LKd^{2}p^{2}b+Vdp_{o}b)$ scales with rank factor $p$. When using LoSiA-Pro without Gradient Checkpointing, only the activations corresponding to the input neurons need to be stored, making it the sole PEFT approach capable of mitigating this class of memory bottlenecks.
Table 14: Comparison of Memory Consumptions. Cells in green highlight the components that may notably lower than other methods, while in red highlight the components that may cause relatively large memory consumption in high-rank.
@@
Table 15: Details of Trainable Parameters for LoSiA under Different Hyperparameter Configurations on LLaMA-2 7B.
@@ A.4.2 Latency Measurement
While demonstrating superior performance among existing baselines, LoSiA reduces training latency by 14.1% compared to LoRA, 55.8% compared to DoRA. The acceleration is mainly due to the elimination of low-rank matrix multiplication. Specifically, during backward propagation with Gradient Checkpointing, the production of low-rank matrices introduces significant overhead for activation recomputation and gradient calculation. Note that LoRA can avoid gradient calculations on backbone weights, but this requires specialized implementations and still introduces a large coefficient of computational complexity.
@@
A.4.3 Algorithm
The core of LoSiA is summarized in Algorithm 2. The model is partitioned into weight groups (in this paper, simply the decoder layers); each group is assigned its own LoSiA optimizer that receives the total number of groups ($L$) and related meta-information, and LoSiA automatically performs the weight updates during the backward pass.
Algorithm 2 Pseudo Code of LoSiA
1:Weight matrix $W\in\mathbb{R}^{n\times m}$ lying in decoder layer (weight groups) $l$; Total decoder layer (weight groups) $L$; EMA ratio $\beta_{1},\beta_{2}$; Adam decay rates $\beta^{\prime}_{1},\beta^{\prime}_{2}$; Rank factor $p$; Time slot $T$;
2:
3:Initialize scales of the core subnet $n_{p}\leftarrow\lfloor np\rfloor,m_{p}\leftarrow\lfloor mp\rfloor$
4:Initialize first- and second-order momentum $M_{0}\leftarrow 0_{n_{p}\times m_{p}}$, $V_{0}\leftarrow 0_{n_{p}\times m_{p}}$
5:Initialize selected neurons randomly $\rho\leftarrow\text{random}([1\ldots n],n_{p})$, $\gamma\leftarrow\text{random}([1\ldots m],m_{p})$
6:Initialize training step $t\leftarrow 1$
7:
8:repeat
9:  $t^{\prime}\leftarrow(t-1)\mod(TL)$
10:  if $\lfloor\frac{t^{\prime}}{T}\rfloor=l-1$ then $\triangleright$ {sensitivity-based parameter importance estimation}
11:   $I\leftarrow W\cdot\nabla W$
12:   $I\leftarrow|I-\frac{1}{2}I^{2}|$ $\triangleright$ {calculation of importance score by Eq.3}
13:
14:   if $t$ is the first step of time slot $T$ then
15:     $\overline{I}_{t-1}\leftarrow 0_{n\times m}$, $\overline{U}_{t-1}\leftarrow 0_{n\times m}$
16:   end if
[…]
18:   $\overline{I}_{t}\leftarrow\beta_{1}\overline{I}_{t-1}+(1-\beta_{1})\overline{I}_{t}$
19:   $\overline{U}_{t}\leftarrow\beta_{2}\overline{U}_{t-1}+(1-\beta_{2})\overline{U}_{t}$
20:  end if
21:  WEIGHTS OPTIMIZATION BY Adam
22:   ($\nabla W_{\rho,\gamma}\leftarrow\tilde{L}_{S}\tilde{R}_{S}$) $\triangleright$ {partial activation $\tilde{L}_{S}$ manually saved during forward in LoSiA-Pro}
23:   $G\leftarrow\nabla W_{\rho,\gamma}$ $\triangleright$ {obtain subnet gradient by indices selection}
24:   $M_{t}\leftarrow\beta^{\prime}_{1}M_{t-1}+(1-\beta^{\prime}_{1})\cdot G$
25:   $V_{t}\leftarrow\beta^{\prime}_{2}V_{t-1}+(1-\beta^{\prime}_{2})\cdot G^{2}$
26:   $M_{t}\leftarrow M_{t}/(1-\beta^{\prime}_{1})$, $V_{t}\leftarrow V_{t}/(1-\beta^{\prime}_{2})$
27:   $N_{t}\leftarrow M_{t}/(\sqrt{V_{t}}+\epsilon)$
28:
29:   $\eta\leftarrow\overline{lr}(t)$ $\triangleright$ {calculate learning rate using rewarming scheduler in Section 3.3}
30:   $W_{t}\leftarrow W_{t-1}-\eta N_{t}$
31:
32:
33:  if $t^{\prime}\mod T=0$ and $\frac{t^{\prime}}{T}=l$ then
34:   $s_{W}\leftarrow\overline{I}_{t-1}\cdot\overline{U}_{t-1}$
35:   for localization algorithm $\mathcal{A}_{i}$ do
36:     $\rho_{i},\gamma_{i}\leftarrow\mathcal{A}_{i}(s_{W},p)$
37:   end for
38:   $k=\arg\max_{i}S(s_{W_{\rho_{i},\gamma_{i}}})$ $\triangleright$ {choose the core subnet with the highest importance}
39:   $\rho,\gamma\leftarrow\rho_{k},\gamma_{k}$
40:   $\text{delete}\ \overline{I},\overline{U}$ $\triangleright$ {delete useless tensors after importance calculation terminate}
41:   $M_{t}\leftarrow 0_{n_{p}\times m_{p}},V_{t}\leftarrow 0_{n_{p}\times m_{p}}$
42:  end if
43:
44:  $t\leftarrow t+1$
45:until training finishes
46:return $W$